# Distributed games with jumps: an $\alpha$-potential game Approach

### v1

@@ 4. Optimize $\alpha$-Potential Function for $\alpha$-NE
$\inf_{{\textbf{{u}}}\in\mathcal{A}}\Phi({\textbf{{u}}}),\quad\Phi({\textbf{{u}}})=\mathbb{E}$ $\left[\int_{0}^{T}F(t,{\textbf{X}}^{{\textbf{{u}}}}_{t},{\textbf{Y}}^{{\textbf{{u}}}}_{t},{\textbf{{u}}}_{t}){\mathrm{d}}t+G({\textbf{X}}^{{\textbf{{u}}}}_{T},{\textbf{Y}}^{{\textbf{{u}}}}_{T})\right],$ (4.1)
@@
${\mathrm{d}}{\textbf{X}}_{t}&=b(t){\textbf{{u}}}_{t}{\mathrm{d}}t+\sigma(t){\mathrm{d}}W_{t}+\sum_{j=1}^{m}\int_{{\mathbb{R}}_{0}^{p}}{\gamma}_{j}(t,z)\widetilde{\eta}_{j}({\mathrm{d}}t,{\mathrm{d}}z),\quad{\textbf{X}}_{0}=\boldsymbol{\xi},\\ {\mathrm{d}}{\textbf{Y}}_{t}&=b(t){\textbf{{u}}}_{t}{\mathrm{d}}t,\quad{\textbf{Y}}_{0}=0,\end{aligned}\right.$ (4.2)
@@ 4.1. Verification theorem
The minimizer of (4.1) can be constructed by standard verification results. To see it, let $\mathcal{C}^{1,2,1}([0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN})$ be the space of functions $\phi=\phi(t,x,y):[0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN}\to{\mathbb{R}}$ such that $\partial_{t}\phi$, $\partial_{x}\phi$, $\partial^{2}_{xx}\phi$, and $\partial_{y}\phi$ exist and are continuous. For all $a\in A$ and $\phi\in\mathcal{C}^{1,2,1}([0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN})$, define the operator $\mathbb{L}^{a}\phi:[0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN}{{\rightarrow}}{\mathbb{R}}$ by
@@
[…]
@@
$\hat{a}(t,x,y)=\arg\min_{a\in A}H(t,x,y,v,a),\quad(t,x,y)\in[0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN},$ (4.4)
@@
${\mathrm{d}}{\textbf{X}}_{t}&=b(t)\hat{a}(t,{\textbf{X}}_{t},{\textbf{Y}}_{t}){\mathrm{d}}t+\sigma(t){\mathrm{d}}W_{t}+\sum_{j=1}^{m}\int_{{\mathbb{R}}_{0}^{p}}{\gamma}_{j}(t,z)\widetilde{\eta}_{j}({\mathrm{d}}t,{\mathrm{d}}z),\quad{\textbf{X}}_{0}=\boldsymbol{\xi},\\ {\mathrm{d}}{\textbf{Y}}_{t}&=b(t)\hat{a}(t,{\textbf{X}}_{t},{\textbf{Y}}_{t}){\mathrm{d}}t,\quad{\textbf{Y}}_{0}=0,\end{aligned}\right.$ (4.5)
has a square integrable strong solution $(\hat{{\textbf{X}}},\hat{{\textbf{Y}}})$ and that the control $\hat{\textbf{{u}}}_{t}\coloneqq\hat{a}(t,\hat{{\textbf{X}}}_{t},\hat{{\textbf{Y}}}_{t})$, $t\in[0,T]$, is in $\mathcal{H}^{2}({\mathbb{R}}^{kN})$. Then $v(0,\xi)=\inf_{{\textbf{{u}}}\in\mathcal{A}}\Phi({\textbf{{u}}})$, and $\hat{\textbf{{u}}}$ is an optimal control of (4.1) and an $\alpha$-NE of the distributed game $\mathcal{G}$, with $\alpha$ given in (3.7).
@@
The proof of Theorem 4.1 follows from standard verification arguments for classical stochastic control problems (see, e.g., [46, Chapter 5]). The first step is to show that $\Phi({\textbf{{u}}})\geq v(0,\xi,0)$ for all ${\textbf{{u}}}\in\mathcal{A}$, by applying Itô’s formula for jump-diffusion processes to the function $t\mapsto v(t,{\textbf{X}}^{{\textbf{{u}}}}_{t},{\textbf{Y}}^{{\textbf{{u}}}}_{t})$ and using the HJB equation (4.3) satisfied by $v$. The second step is to show that $\Phi(\hat{{\textbf{{u}}}})=v(0,\xi,0)$ due to the definition (4.4) of $\hat{a}$, which implies the optimality of $\hat{\textbf{{u}}}$.
@@ 4.2. Viscosity characterization
$V(t,x,y)\coloneqq\inf_{{\textbf{{u}}}\in\mathcal{A}}\mathbb{E}$ $\left[\int_{t}^{T}F(s,{\textbf{X}}^{{\textbf{{u}}}}_{t},{\textbf{Y}}^{{\textbf{{u}}}}_{s},{\textbf{{u}}}_{s}){\mathrm{d}}s+G({\textbf{X}}^{{\textbf{{u}}}}_{T},{\textbf{Y}}^{{\textbf{{u}}}}_{T})\bigg{|}X^{{\textbf{{u}}}}_{t}=x,Y^{{\textbf{{u}}}}_{t}=y\right].$ (4.6)
@@
Now we identify the value function $V$ defined by (4.6) as the unique viscosity solution to (4.3). As in [21, Definition 2.1], we say a function $v:[0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN}\to{\mathbb{R}}$ is a viscosity subsolution (resp. supersolution) of (4.3) if $v$ is upper semicontinuous (resp. lower semicontinuous) and for every $(t_{0},x_{0},y_{0})\in[0,T)\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN}$ and $\phi\in C^{1,2,1}([0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN})$ such that $\phi-v$ attains its minimum (resp. maximum) at $(t_{0},x_{0},y_{0})$,
[…]
@@
Theorem 4.2 is a fairly standard result; however, as most existing references focus on the case where the cost functions $F$ and $G$ are globally Lipschitz continuous (see, e.g., [35, 12, 4]), we include a brief sketch of the proof below for completeness. Under (H.2), the functions $F$ and $G$ defined in (3.6) are continuous and it holds for some $C\geq 0$ that for all $t\in[0,T]$ and $(x,y),(x^{\prime},y^{\prime})\in{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN}$, $|F(t,x,y)|+|G(x,y)|\leq C(1+|x|^{2}+|y|^{2})$ and
$|F(t,x,y)-F(t,x^{\prime},y^{\prime})|+|G(x,y)-G(x^{\prime},y^{\prime})|\leq C(1+|x|+|y|+|x^{\prime}|+|y^{\prime}|)(|x-x^{\prime}|+|y-y^{\prime}|).$
[…]
@@ 5. Policy Gradient Algorithm for $\alpha$-NE
The algorithm begins by approximating the NE policy given in Theorem 4.1 using a sufficiently expressive parametric family (e.g., a family of deep neural networks). Specifically, we consider a family of policy profiles $\phi_{\theta}:[0,T]\times\mathbb{R}^{dN}\times\mathbb{R}^{dN}\to A$ with weights $\theta\in\mathbb{R}^{L}$, and consider for each $\theta\in{\mathbb{R}}^{L}$,
$\Phi(\theta)\coloneqq\mathbb{E}$ $\left[\int_{0}^{T}F(t,{\textbf{X}}^{\theta}_{t},{\textbf{Y}}^{\theta}_{t},\phi_{\theta}(t,\mathbf{X}^{\theta}_{t},\mathbf{Y}^{\theta}_{t})){\mathrm{d}}t+G({\textbf{X}}^{\theta}_{T},{\textbf{Y}}^{\theta}_{T})\right],$ (5.1)
@@
${\mathrm{d}}{\textbf{X}}_{t}&=b(t)\phi_{\theta}(t,{\textbf{X}}_{t},{\textbf{Y}}_{t}){\mathrm{d}}t+\sigma(t){\mathrm{d}}W_{t}+\sum_{j=1}^{m}\int_{{\mathbb{R}}_{0}^{p}}{\gamma}_{j}(t,z)\widetilde{\eta}_{j}({\mathrm{d}}t,{\mathrm{d}}z),\quad{\textbf{X}}_{0}=\boldsymbol{\xi},\\ {\mathrm{d}}{\textbf{Y}}_{t}&=b(t)\phi_{\theta}(t,{\textbf{X}}_{t},{\textbf{Y}}_{t}){\mathrm{d}}t,\quad{\textbf{Y}}_{0}=0,\end{aligned}\right.$ (5.2)
@@
$X^{\theta}_{i,t_{\ell+1}}$ $=X^{\theta}_{i,t_{\ell}}+b_{i}(t_{\ell})\phi_{\theta}(t_{\ell},{\textbf{X}}^{\theta}_{t_{\ell}},{\textbf{Y}}^{\theta}_{t_{\ell}})\Delta_{\ell}+\sigma_{i}(t_{\ell})\Delta W_{\ell}$ (5.3)
@@
where $\Delta_{\ell}\coloneqq t_{\ell+1}-t_{\ell}$, $\Delta W_{\ell}\coloneqq W_{t_{\ell+1}}-W_{t_{\ell}}$, $N_{j,\ell}$ denotes the number of jumps of the $j$-th Poisson random measure occurring over the time interval $[0,t_{\ell}]$, and $z_{k}$ is the size of the $k$-th jump sampled from the distribution $\nu/\nu({\mathbb{R}}^{p}_{0})$. Let $({\textbf{X}}^{\theta,(m)},{\textbf{Y}}^{\theta,(m)})_{m=1}^{M}$, $M\in{\mathbb{N}}$, be independent trajectories of (5.3) with policy $\phi_{\theta}$, and define the following empirical approximation of (5.1)
$\Phi_{M}(\theta)\coloneqq\frac{1}{M}\sum_{m=1}^{M}\left[\sum_{\ell=0}^{P-1}F\left({\textbf{X}}^{\theta,(m)}_{t_{\ell}},{\textbf{Y}}^{\theta,(m)}_{t_{\ell}},\phi_{\theta}\left(t_{\ell},{\textbf{X}}^{\theta,(m)}_{t_{\ell}},{\textbf{Y}}^{\theta,(m)}_{t_{\ell}}\right)\right)\Delta_{\ell}+G\left({\textbf{X}}^{\theta,(m)}_{t_{P}},{\textbf{Y}}^{\theta,(m)}_{t_{P}}\right)\right].$ (5.4)

### v2

@@ 4. Optimize $\alpha$-Potential Function for $\alpha$-NE
(4.1) $\inf_{{\textbf{{u}}}\in\mathcal{A}}\Phi({\textbf{{u}}}),\quad\Phi({\textbf{{u}}})=\mathbb{E}$ $\left[\int_{0}^{T}F(t,{\textbf{X}}^{{\textbf{{u}}}}_{t},{\textbf{Y}}^{{\textbf{{u}}}}_{t},{\textbf{{u}}}_{t}){\mathrm{d}}t+G({\textbf{X}}^{{\textbf{{u}}}}_{T},{\textbf{Y}}^{{\textbf{{u}}}}_{T})\right],$
@@
(4.2) $\left\{\begin{aligned} {\mathrm{d}}{\textbf{X}}_{t}&=b(t){\textbf{{u}}}_{t}{\mathrm{d}}t+\sigma(t){\mathrm{d}}W_{t}+\sum_{j=1}^{m}\int_{{\mathbb{R}}_{0}^{p}}{\gamma}_{j}(t,z)\widetilde{\eta}_{j}({\mathrm{d}}t,{\mathrm{d}}z),\quad{\textbf{X}}_{0}=\boldsymbol{\xi},\\ {\mathrm{d}}{\textbf{Y}}_{t}&=b(t){\textbf{{u}}}_{t}{\mathrm{d}}t,\quad{\textbf{Y}}_{0}=0,\end{aligned}\right$.
@@ 4.1. Verification theorem
The minimizer of (4.1) can be constructed by standard verification results. To see it, let $\mathcal{C}^{1,2,1}([0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN})$ be the space of functions $\phi=\phi(t,x,y):[0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN}\to{\mathbb{R}}$ such that $\partial_{t}\phi$, $\partial_{x}\phi$, $\partial^{2}_{xx}\phi$, and $\partial_{y}\phi$ exist and are continuous and there exists $C\geq 0$ such that for all $(t,x,y)\in[0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN}$, $|\phi(t,x,y)|+|\partial_{t}\phi(t,x,y)|\leq C(1+|x|^{2}+|y|^{2})$, $|\partial_{x}\phi(t,x,y)|+|\partial_{y}\phi(t,x,y)|\leq C(|1+|x|+|y|)$, and $|\partial^{2}_{x}\phi(t,x,y)|\leq C$. For all $a\in A$ and $\phi\in\mathcal{C}^{1,2,1}([0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN})$, define the operator $\mathbb{L}^{a}\phi:[0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN}{{\rightarrow}}{\mathbb{R}}$ by
@@
(4.3) $\left\{\begin{aligned} &\partial_{t}w(t,x,y)+\inf_{a\in A}H(t,x,y,w,a)=0,\quad(t,x,y)\in[0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN},\\ &w(T,x,y)=G(x,y),\quad(x,y)\in{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN},\end{aligned}\right$.
@@
(4.4) $\hat{a}(t,x,y)=\arg\min_{a\in A}H(t,x,y,v,a),\quad(t,x,y)\in[0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN},$
@@
(4.5) $\left\{\begin{aligned} {\mathrm{d}}{\textbf{X}}_{t}&=b(t)\hat{a}(t,{\textbf{X}}_{t},{\textbf{Y}}_{t}){\mathrm{d}}t+\sigma(t){\mathrm{d}}W_{t}+\sum_{j=1}^{m}\int_{{\mathbb{R}}_{0}^{p}}{\gamma}_{j}(t,z)\widetilde{\eta}_{j}({\mathrm{d}}t,{\mathrm{d}}z),\quad{\textbf{X}}_{0}=\boldsymbol{\xi},\\ {\mathrm{d}}{\textbf{Y}}_{t}&=b(t)\hat{a}(t,{\textbf{X}}_{t},{\textbf{Y}}_{t}){\mathrm{d}}t,\quad{\textbf{Y}}_{0}=0,\end{aligned}\right$.
@@
Theorem 4.1 indicates that under sufficient regularity conditions, an $\alpha$-NE for the game $\mathcal{G}$ can be obtained by minimizing the $\alpha$-potential function $\Phi$ in (4.1) over feedback controls of the form ${\textbf{{u}}}_{t}=\phi(t,{\textbf{X}}_{t},{\textbf{Y}}_{t})$, where $\phi:[0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN}\to A$ is a sufficiently regular policy profile, and $({\textbf{X}},{\textbf{Y}})$ satisfies (4.5) with $\hat{a}$ replaced by $\phi$. This result provides the theoretical foundation for the policy gradient algorithm presented in Section 5.
The proof of Theorem 4.1 follows from standard verification arguments for classical stochastic control problems (see, e.g., [56, Chapter 5]). The first step is to show that $\Phi({\textbf{{u}}})\geq v(0,\xi,0)$ for all ${\textbf{{u}}}\in\mathcal{A}$, by applying Itô’s formula for jump-diffusion processes to the function $t\mapsto v(t,{\textbf{X}}^{{\textbf{{u}}}}_{t},{\textbf{Y}}^{{\textbf{{u}}}}_{t})$ and using the HJB equation (4.3) satisfied by $v$. The second step is to show that $\Phi(\hat{{\textbf{{u}}}})=v(0,\xi,0)$ due to the definition (4.4) of $\hat{a}$, which implies the optimality of $\hat{\textbf{{u}}}$.
@@ 4.2. Viscosity characterization
(4.6) $V(t,x,y)\coloneqq\inf_{{\textbf{{u}}}\in\mathcal{A}}\mathbb{E}$ $\left[\int_{t}^{T}F(s,{\textbf{X}}^{{\textbf{{u}}}}_{t},{\textbf{Y}}^{{\textbf{{u}}}}_{s},{\textbf{{u}}}_{s}){\mathrm{d}}s+G({\textbf{X}}^{{\textbf{{u}}}}_{T},{\textbf{Y}}^{{\textbf{{u}}}}_{T})\bigg|X^{{\textbf{{u}}}}_{t}=x,Y^{{\textbf{{u}}}}_{t}=y\right].$
@@
Now we identify the value function $V$ defined by (4.6) as the unique viscosity solution to (4.3). We first recall the definition of viscosity solutions as given in [25, Definition 2.1]:
Definition 4.2.
A function $v:[0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN}\to{\mathbb{R}}$ is called a viscosity subsolution (resp. supersolution) of (4.3) if $v$ is upper semicontinuous (resp. lower semicontinuous) and for every $(t_{0},x_{0},y_{0})\in[0,T)\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN}$ and $\phi\in C^{1,2,1}([0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN})$ such that $\phi-v$ attains its minimum (resp. maximum) at $(t_{0},x_{0},y_{0})$,
$\partial_{t}\phi(t_{0},x_{0},y_{0})+\inf_{a\in A}H(t_{0},x_{0},y_{0},\phi,a)\geq 0\quad\textnormal{(resp.~$\leq 0)$}.$
If $v$ is both a viscosity subsolution and a viscosity supersolution of (4.3), then $v$ is called a viscosity solution.
@@
Proof.
The proof follows from standard arguments in viscosity solution theory, adapting results for control problems with Lipschitz continuous costs (see, e.g., [42, 12]) to the present setting with locally Lipschitz continuous costs. Indeed, under (H.2), the functions $F$ and $G$ defined in (3.6) are continuous, and there exists $C\geq 0$ such that for all $t\in[0,T]$, $(x,y),(x^{\prime},y^{\prime})\in{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN}$ and $a\in A$,
(4.7) $|F(t,x,y,a)|+|G(x,y)|\leq C(1+|x|^{2}+|y|^{2}),$
$|F(t,x,y,a)-F(t,x^{\prime},y^{\prime},a)|+|G(x,y)-G(x^{\prime},y^{\prime})|$
(4.8) $\quad\leq C(1+|x|+|y|+|x^{\prime}|+|y^{\prime}|)(|x-x^{\prime}|+|y-y^{\prime}|).$
To see it, a direct computation and the boundedness of $\partial_{xx}g_{i}$ imply that
$|\partial_{x_{i}}G(x,y)|$ $=\left|\sum_{j=1}^{N}\int_{0}^{1}(\partial_{x_{i}x_{j}}g_{j})(x-(1-r)y)y_{j}{\mathrm{d}}r\right|\leq C|y|,$
$|\partial_{y_{i}}G(x,y)|$ $=\left|\int_{0}^{1}(\partial_{x_{i}}g_{i})(x-(1-r)y){\mathrm{d}}r-\sum_{j=1}^{N}\int_{0}^{1}(1-r)(\partial_{x_{i}x_{j}}g_{j})(x-(1-r)y)y_{j}{\mathrm{d}}r\right|$
$\leq C(1+|x|+|y|),$
for a constant $C\geq 0$ independent of $(t,x,y)$. This along with the mean value theorem proves the local Lipschitz continuity and the quadratic growth of $G$ given in (4.7) and (4.8). Applying a similar argument to $F$ establishes the analogous regularity of $F$.
By (4.7) and [12, Lemmas 4.2], the value function $V$ has at most quadratic growth in $(x,y)$. Define the upper semicontinuous envelope $V^{*}$ and the lower semicontinuous envelope $V_{*}$ of $V$ such that for all $(t,x,y)\in[0,T]\times{\mathbb{R}}^{dN}\times{\mathbb{R}}^{dN}$,
$V^{*}(t,x,y)=\limsup_{(t^{\prime},x^{\prime},y^{\prime})\to(t,x,y)}V(t,x,y),\quad V_{*}(t,x,y)=\liminf_{(t^{\prime},x^{\prime},y^{\prime})\to(t,x,y)}V(t,x,y).$
The quadratic growth of $V$ implies that $V^{*}$ and $V_{*}$ have at most quadratic growth. As shown in [12, Theorem 5.2], by the weak dynamic programming principle, $V^{*}$ and $V_{*}$ are viscosity subsolution and supersolution of (4.3), respectively. The local Lipschitz continuity of $F$ and $G$ in (4.8) and the strong comparison principle in [25, Theorem 4.3] imply that $V^{*}\leq V_{*}$, which along with the fact that $V^{*}\geq V\geq V_{*}$ yields that $V$ is the unique continuous viscosity solution. ∎
@@ 5. Policy Gradient Algorithm for $\alpha$-NE
The algorithm begins by approximating the NE policy given in Theorem 4.1 using a sufficiently expressive parametric family (e.g., a family of deep neural networks) [21, 14]. Specifically, we consider a family of policy profiles $\phi_{\theta}:[0,T]\times\mathbb{R}^{dN}\times\mathbb{R}^{dN}\to A$ with weights $\theta\in\mathbb{R}^{L}$, and consider for each $\theta\in{\mathbb{R}}^{L}$,
(5.1) $\Phi(\theta)\coloneqq\mathbb{E}$ $\left[\int_{0}^{T}F(t,{\textbf{X}}^{\theta}_{t},{\textbf{Y}}^{\theta}_{t},\phi_{\theta}(t,\mathbf{X}^{\theta}_{t},\mathbf{Y}^{\theta}_{t})){\mathrm{d}}t+G({\textbf{X}}^{\theta}_{T},{\textbf{Y}}^{\theta}_{T})\right],$
@@
(5.2) $\left\{\begin{aligned} {\mathrm{d}}{\textbf{X}}_{t}&=b(t)\phi_{\theta}(t,{\textbf{X}}_{t},{\textbf{Y}}_{t}){\mathrm{d}}t+\sigma(t){\mathrm{d}}W_{t}+\sum_{j=1}^{m}\int_{{\mathbb{R}}_{0}^{p}}{\gamma}_{j}(t,z)\widetilde{\eta}_{j}({\mathrm{d}}t,{\mathrm{d}}z),\quad{\textbf{X}}_{0}=\boldsymbol{\xi},\\ {\mathrm{d}}{\textbf{Y}}_{t}&=b(t)\phi_{\theta}(t,{\textbf{X}}_{t},{\textbf{Y}}_{t}){\mathrm{d}}t,\quad{\textbf{Y}}_{0}=0,\end{aligned}\right$.
@@
(5.3) $X^{\theta}_{i,t_{\ell+1}}$ $=X^{\theta}_{i,t_{\ell}}+b_{i}(t_{\ell})\phi_{\theta}(t_{\ell},{\textbf{X}}^{\theta}_{t_{\ell}},{\textbf{Y}}^{\theta}_{t_{\ell}})\Delta_{\ell}+\sigma_{i}(t_{\ell})\Delta W_{\ell}$
@@
where $\Delta_{\ell}\coloneqq t_{\ell+1}-t_{\ell}$, $\Delta W_{\ell}\coloneqq W_{t_{\ell+1}}-W_{t_{\ell}}$, $N_{j,\ell}$ denotes the number of jumps of the $j$-th Poisson random measure occurring over the time interval $0,t_{\ell}]$, and $z_{k}$ is the size of the $k$-th jump sampled from the distribution $\nu/\nu({\mathbb{R}}^{p}_{0})$. Let $({\textbf{X}}^{\theta,(m)},{\textbf{Y}}^{\theta,(m)})_{m=1}^{M}$, $M\in{\mathbb{N}}$, be independent trajectories of (5.3) with policy $\phi_{\theta}$, and define the following empirical approximation of (5.1)
(5.4) $\Phi_{M}(\theta)\coloneqq\frac{1}{M}\sum_{m=1}^{M}\left[\sum_{\ell=0}^{P-1}F\left({\textbf{X}}^{\theta,(m)}_{t_{\ell}},{\textbf{Y}}^{\theta,(m)}_{t_{\ell}},\phi_{\theta}\left(t_{\ell},{\textbf{X}}^{\theta,(m)}_{t_{\ell}},{\textbf{Y}}^{\theta,(m)}_{t_{\ell}}\right)\right)\Delta_{\ell}+G\left({\textbf{X}}^{\theta,(m)}_{t_{P}},{\textbf{Y}}^{\theta,(m)}_{t_{P}}\right)\right$].